# MARFT: Multi-Agent Reinforcement Fine-Tuning

### v2

@@ 3.1 Problem Statement and Flex-POMDP
To enhance problem formulation and accommodate MARFT, we propose Flexible Partially Observable Markov Decision Process (Flex-POMDP), denoted as $\langle\mathcal{V},\mathcal{N},{\mathcal{S}},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R},\gamma,\mathcal{D}\rangle$, where $\mathcal{N}=\{1,...,n\}$ is the agent composition of a LaMAS with some organization, i.e., some order constraint among agents. $\mathcal{V},\mathcal{S},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R}\ \text{and}\ \gamma$ can be borrowed from Section 2.3.1. A key contribution of the Flex-POMDP framework is the introduction of a dependency function $\mathcal{D}\colon\bm{\mathcal{A}}\times\bm{\mathcal{A}}\rightarrow\{0,1\}$, which explicitly models inter-agent dependencies at the level of action execution or outcome. Specifically, $\mathcal{D}(a^{i},a^{j})=1$ indicates that the action of agent $j$ is conditionally dependent on the action of agent $i$. As a result, the decision-making process of agent $j$ is influenced not only by its own observation but also by the actions of all agents $k$ such that $k\in\{l|\mathcal{D}(a^{l},a^{j})=1\}$. These dependencies can be integrated into agent $j$’s input via concatenation or other fusion methods, thereby forming an enriched "rollout observation". Crucially, the dependency function $\mathcal{D}$ is allowed to vary across timesteps, potentially governed by an orchestration agent or a dynamic routing mechanism. This flexibility captures the evolving coordination structures in LaMAS and embodies the "flexible" nature of the proposed framework. Notably, when $\mathcal{D}(a^{i},a^{j})=0$ for all $i,j$, the Flex-POMDP reduces to the standard DEC-POMDP, thereby demonstrating the generality of the formulation.
@@ 3.2 Categorizations of LaMAS in MARFT
Parameter sharing. Parameter sharing is a critical consideration in multi-agent systems, particularly in traditional MARL. However, in LaMAS, parameter differentiation strategies introduce additional complexity. Agents may either share parameters or maintain distinct configurations, each with unique implications for optimization.
@@
Execution synchronicity. Execution synchronicity represents another critical dimension for classifying LaMAS. Unlike traditional cooperative games in multi-agent systems, where agents typically act synchronously, LaMAS often operates under asynchronous or hybrid execution paradigms. This distinction arises from the inherent dependencies in agentic tasks, where execution outcomes frequently rely on partial or sequential results.
@@
Update simultaneity. In addition to execution synchronicity during inference, the choice of update scheme during training further distinguishes LaMAS configurations. Specifically, LaMAS can be categorized into two types based on how agent updates are orchestrated.
@@ 3.3 Differences and Resulting Challenges
Table 2: Differences between MARFT and traditional MARL.
@@
Action Asynchronicity. One of the most significant differences between conventional MARL and MARFT is the nature of agent actions. In traditional cooperative MARL, agents typically act simultaneously. However, in MARFT, agents often act asynchronously. In some cases, the actions of certain agents may even depend on the outcomes of other agents’ actions. For example, in a collaborative coding assistant system, one LLM agent generates a function prototype, and another agent asynchronously refines it based on the first agent’s output, demonstrating both asynchronicity and result dependency. As a result, conventional MARL methods, which assume synchronous actions, may not be directly applicable or effective in LLM-based multi-agent systems.
@@
Utilities. Unlike agents in typical MARL problems, LLMs are initially designed for language processing rather than specific agentic tasks. However, the value- or reward-guided nature of RL means that when using value-based RL algorithms to optimize or extract policies, the derived optimal policy often prioritizes actions that maximize the value function. This can lead to policies that generate high-reward actions or tokens but are incomprehensible to humans. Though some attempts to extract a policy together with entropy regularization have been made to improve its agentic intelligence without harming the text capability (Wen et al., 2024), it is still a point that requires additional attention when we try to implement value-based methods to optimize LLMs.
@@
Characteristic Profiles. The transition from single-agent to multi-agent systems in typical MARL benchmarks, such as Multi-Agent MuJoCo, often involves splitting a single agent into multiple components without additional modifications. In contrast, LLMs require a more nuanced approach. When decomposing tasks into orthogonal sub-tasks, each LLM agent needs a profile to define its role and capabilities, enabling it to generate actions consistent with its assigned character. This profile can be human-specified or learned by the agent through natural evolution. Consequently, the joint observation space in MARFT is augmented with profiles, taking the form [agent profile + env observation (+ agent-specific guidance)]. This modification is crucial when designing a MARL training framework.
@@
Heterogeneity. In conventional MARL benchmarks, heterogeneity is typically characterized by differences in agent structures and non-parameter-sharing schemes. However, LLM-based systems introduce a higher level of complexity. First, LLMs themselves can vary significantly in terms of model structures, parameters, input and output formats (e.g. , LLMs vs. Vision-Language Models), and vocabularies. Second, agents may have access to different external tools or devices, reflecting real-world scenarios. For instance, in designing a multi-agent system for a mobile operating system, some agents may have access to both local and remote search engines, while others may rely on proprietary databases or tools from other companies. This heterogeneity complicates training, making it more difficult and unstable.
@@
System Organization. Traditional MARL tasks are often set in static simulated environments. In contrast, LLM-based multi-agent systems are designed for real-life agentic tasks with higher uncertainty. These tasks can be decomposed in various ways, leading to different multi-agent systems with distinct populations and roles. Moreover, agents may exhibit sequential dependencies and contextual relationships when solving sub-tasks. For example, one agent’s action may depend on the outcome of another agent’s action. This dynamic organization can be either human-designed or agent-explored (e.g. , through training), adding another layer of complexity to the design process.
@@
Optimization Space. If we take one LLM generation as an action, the observation space is exponentially vast as $o\in\mathcal{O}\subseteq\mathcal{V}^{L}$, whose complexity is $O(\lvert\mathcal{V}\rvert^{L})$, where $\mathcal{V}$ is the vocabulary of the acting LLM and $N$ is the token length of the generated action, and both of them vary with different acting LLMs. It makes the value function hard to converge with stability. If we take one token as an action, the complexity can be reduced to $O(\lvert\mathcal{V}\rvert\times L)$, but it will become a task with super sparse reward outcome so that the value function should also be meticulously learned, as $L$ can be extremely large. This is also reflected in VAPO and is mitigated by value pretraining (Yue et al., 2025). Fortunately, not all LLM-related tasks face such large optimization spaces. For instance, in embodied AI with LLMs, predefined actions (e.g. , "go to the kitchen," "grab a coffee") are often provided, significantly narrowing the action space (Carta et al., 2023; Tan et al., 2024).
These differences collectively make MARFT problems more complex than traditional MARL problems. In the following section, we will explore some mature MARFT algorithms, building on existing methods.
@@ 4 Multi-Agent Reinforcement Fine-Tuning
In this section, we will dive into the inheriting methods for MARFT, and these methods are rooted in the most substantial RL policy optimization methods, such as TRPO, PPO, because they have been proven to give a monotonous policy improvement guarantee.
@@ 4.1.1 Action-level Fine-Tuning
Action-level MARFT. To tackle the difficulties listed above, naive MARFT inherits the SM-style modeling to unify the dynamic and undefined LaMAS decision processes perfectly and trust-region learning to guarantee monotonic improvement on the performance of LaMAS.
@@
Given any organization or task-solving workflow of a LaMAS, via the action advantage decomposition theorem, MARFT re-models it into a sequential decision-making process. And then during training, MARFT incorporates the transformer-like problem-solving style and trust-region learning. The encoder here is to form the local roll-out observations with agents’ profiles or other information, and the decoder here is to combine the roll-out observations with the actions of the predecessors and generate the action of the currently acting agent. Since actions are taken sequentially as $a^{i}\sim p_{\pi^{i}}(a^{i}|\hat{o}^{i},a^{i-1})$ whose log probability is computed by (20), which perfectly gives the probabilities MAT training needs, the optimization objective demonstrated as (19) can also be borrowed seamlessly. It is worth noting that, in naive MARFT, an agent optimizes in a trust region conditioned on the predecessors’ actions as in MAT, so it gives a monotonic improvement guarantee like the sequential update scheme in HAPPO.
$\text{log}P_{\pi^{i}}\left(a^{i}|\hat{o}^{i},\bm{a}^{i-1}\right)$ $=\text{log}P_{\pi^{i}}\left(w^{i,1},\ldots,w^{i,L^{i}}|\hat{o}^{i},\bm{a}^{i-1}\right)$ (20)
$=\sum_{j=1}^{L^{i}}\text{log}P_{\pi^{i}}\left(w^{i,j}|\hat{o}^{i},\bm{a}^{i-1},w^{i,1},\ldots,w^{i,j-1}\right).$ (21)
@@
Action Normalization. As shown in equation (20), the probability of each token $P\left(w|\cdot\right)$ is less than (or equal to) 1. As a result, longer action tends to have a lower joint probability, even though it is more reasonable than other actions in the task. To address the issue, Tan et al. (2024) proposed token normalization and word normalization by dividing the joint action probability in (20) by a constant related to its token or word length. It gets rid of the strong assumption that all the actions have similar lengths, which in LLMs, can be an evil condition, because in agentic tasks, it is all too common that solutions generated can be either super long with a chain-of-thought or super short with a direct answer.
@@
Agent-by-agent Update. Updating every policy during each training epoch may introduce implicit off-policy issues, which can potentially compromise the monotonic improvement guarantee of PPO for individual agents’ policies. This problem can be exacerbated in LaMAS, where agents often have dependencies on the results of other agents’ actions. Drawing on the approach of A2PO (Wang et al., 2023), an agent-by-agent update scheme can be employed to restore monotonic improvement for each agent and facilitate stable convergence.
@@
Trajectory-level Optimization. For language models, the definition of action from the RL perspective can be dynamic. Direct trajectory-level optimization, such as GRPO with outcome supervision (Shao et al., 2024) and its variant DAPO (Yu et al., 2025), can be viewed as an action-level optimization with non-zero reward only at the end of the trajectory in the MDP or a normal action-level optimization in a single-step variant of the MDP. Consequently, trajectory-level optimization can be seamlessly integrated into MARFT by following the same idea.
@@ 4.1.2 Token-level Fine-Tuning
Token-level MARFT. On top of the credit assignment of each agent on its action, we can also give a more fine-grained credit assignment by naturally treating every token as an action. Before showing the difference in algorithms, we need to define the token-level multi-agent state-action and state value functions as
$Q^{\bm{\pi}}\left(s_{t},\bm{a}_{t}^{1:i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)\coloneqq\mathbb{E}_{p_{\bm{\pi}}\left(\tau|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)}\left[\mathcal{R}^{\gamma}|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right],$ (22)
$V^{\bm{\pi}}\left(s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)}\left[\mathcal{R}^{\gamma}|s_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right],$ (23)
where $\mathcal{R}^{\gamma}=\sum_{k=t}^{\infty}\gamma^{k-t}r\left(s_{k},\bm{a}_{k}\right)$. And then the token-level Bellman backup can be spontaneously derived as
$Q^{\bm{\pi}}(o_{t},a_{t}^{i-1},w_{t}^{i,1:j-1},w_{t}^{i,j})\leftarrow\begin{cases}0+\gamma\max\limits_{w_{t}^{i,j+1}}Q^{\bm{\pi}}(o_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j},w_{t}^{i,j+1})&\mathrm{if~{}}j<|a_{t}^{i}|\\ 0+\gamma\max\limits_{w_{t}^{i+1,1}}Q^{\bm{\pi}}(o_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}{i+1<n}\\ R(o_{t},\bm{a}_{t}^{1:i})+\gamma\max\limits_{w_{t+1}^{1,1}}Q^{\bm{\pi}}(o_{t+1},\bm{a}_{t}^{1:i},w_{t+1}^{1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}i=n\end{cases},$ (24)
$V^{\bm{\pi}}(o_{t},a_{t}^{i-1},w_{t}^{i,1:j})\leftarrow\begin{cases}0+\gamma V^{\bm{\pi}}(o_{t},\bm{a}_{t}^{1:i-1},w_{t}^{i,1:j+1})&\mathrm{if~{}}j<|a_{t}^{i}|\\ 0+\gamma V^{\bm{\pi}}(o_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}{i<n}\\ R(o_{t},\bm{a}_{t}^{1:i})+\gamma V^{\bm{\pi}}(o_{t+1},\bm{a}_{t}^{1:i},\emptyset)&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}i=n\end{cases}.$ (25)
In the context of the naive token-level MARFT approach, each token is treated as an individual action, thereby introducing a new POMDP formulation. This results in a highly sparse reward function, which assigns zero rewards to all intermediate tokens and preceding agents, with the reward being allocated solely to the final token of the entire multi-agent system. Although this sparsity in the reward signal poses challenges for learning the value function, it significantly reduces the optimization complexity from $O\left(\lvert\mathcal{V}\rvert^{L}\right)$ to $O\left(\lvert\mathcal{V}\rvert\times L\right)$, as previously discussed in Section 3.3.

### v3

@@ 3.1 Problem Statement and Flex-POMDP
Figure 2: A detailed illustration of the dynamics of a Flex-POMDP. The dependency function (dashed purple line) can vary across timesteps.
To enhance problem formulation and accommodate MARFT, we propose Flexible Partially Observable Markov Decision Process (Flex-POMDP), illustrated in Figure 2 and denoted as $\langle\mathcal{V},\mathcal{N},{\mathcal{S}},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R},\gamma,\mathcal{D}\rangle$, where $\mathcal{N}=\{1,...,n\}$ is the agent composition of a LaMAS with some organization, i.e., some order constraint among agents. $\mathcal{V},\mathcal{S},\bm{\mathcal{O}},\bm{\mathcal{A}},\mathcal{T},\mathcal{R}\ \text{and}\ \gamma$ can be borrowed from Section 2.3.1. Flex-POMDP extends the DEC-POMDP (Oliehoek & Amato, 2016) by introducing a dependency function $\mathcal{D}\colon\bm{\mathcal{A}}\times\bm{\mathcal{A}}\rightarrow\{0,1\}$, which explicitly models inter-agent dependencies at the level of action execution or outcome. Specifically, $\mathcal{D}(a^{i},a^{j})=1$ indicates that the action of agent $j$ is conditionally dependent on the action of agent $i$. As a result, the decision-making process of agent $j$ is influenced not only by its own observation but also by the actions of all agents $k$ such that $k\in\{l|\mathcal{D}(a^{l},a^{j})=1\}$. These dependencies can be integrated into agent $j$’s input via concatenation or other fusion methods, thereby forming an enriched "rollout observation". Crucially, the dependency function $\mathcal{D}$ can vary across timesteps, potentially governed by an orchestration agent or a dynamic routing mechanism. This flexibility captures the evolving coordination structures in LaMAS and embodies the "flexible" nature of the proposed framework. Notably, when $\mathcal{D}(a^{i},a^{j})=0$ for all $i,j$, Flex-POMDP reduces to the standard DEC-POMDP, thereby demonstrating the generality of the formulation.
@@ 3.2 Categorizations of LaMAS in MARFT
Parameter sharing
Parameter sharing is a critical consideration in multi-agent systems, particularly in traditional MARL. However, in LaMAS, parameter differentiation strategies introduce additional complexity. Agents may either share parameters or maintain distinct configurations, each with unique implications for optimization.
@@
Execution synchronicity
Execution synchronicity represents another critical dimension for classifying LaMAS. Unlike traditional cooperative games in multi-agent systems, where agents typically act synchronously, LaMAS often operates under asynchronous or hybrid execution paradigms. This distinction arises from the inherent dependencies in agentic tasks, where execution outcomes frequently rely on partial or sequential results.
@@
Update simultaneity
In addition to execution synchronicity during inference, the choice of update scheme during training further distinguishes LaMAS configurations. Specifically, LaMAS can be categorized into two types based on how agent updates are orchestrated.
@@ 3.3 Differences and Resulting Challenges
Table 2: Differences between MARFT and traditional MARL.
@@
Action Asynchronicity
@@
One of the most significant differences between conventional MARL and MARFT is the nature of agent actions. In traditional cooperative MARL, agents typically act simultaneously. However, in MARFT, agents often act asynchronously. In some cases, the actions of certain agents may even depend on the outcomes of other agents’ actions. For example, in a collaborative coding assistant system, one LLM agent generates a function prototype, and another agent asynchronously refines it based on the first agent’s output, demonstrating both asynchronicity and result dependency. As a result, conventional MARL methods, which assume synchronous actions, may not be directly applicable or effective in LLM-based multi-agent systems.
Utilities
@@
Unlike agents in typical MARL problems, LLMs are initially designed for language processing rather than specific agentic tasks. However, the value- or reward-guided nature of RL means that when using value-based RL algorithms to optimize or extract policies, the derived optimal policy often prioritizes actions that maximize the value function. This can lead to policies that generate high-reward actions or tokens but are incomprehensible to humans. Though some attempts to extract a policy together with entropy regularization have been made to improve its agentic intelligence without harming the text capability (Wen et al., 2024a), it is still a point that requires additional attention when we try to implement value-based methods to optimize LLMs.
Characteristic Profiles
@@
The transition from single-agent to multi-agent systems in typical MARL benchmarks, such as Multi-Agent MuJoCo, often involves splitting a single agent into multiple components without additional modifications. In contrast, LLMs require a more nuanced approach. When decomposing tasks into orthogonal sub-tasks, each LLM agent needs a profile to define its role and capabilities, enabling it to generate actions consistent with its assigned character. This profile can be human-specified or learned by the agent through natural evolution. Consequently, the joint observation space in MARFT is augmented with profiles, taking the form [agent profile + env observation (+ agent-specific guidance)]. This modification is crucial when designing a MARL training framework.
Heterogeneity
@@
In conventional MARL benchmarks, heterogeneity is typically characterized by differences in agent structures and non-parameter-sharing schemes. However, LLM-based systems introduce a higher level of complexity. First, LLMs themselves can vary significantly in terms of model structures, parameters, input and output formats (e.g. , LLMs vs. Vision-Language Models), and vocabularies. Second, agents may have access to different external tools or devices, reflecting real-world scenarios. For instance, in designing a multi-agent system for a mobile operating system, some agents may have access to both local and remote search engines, while others may rely on proprietary databases or tools from other companies. This heterogeneity complicates training, making it more difficult and unstable.
System Organization
@@
Traditional MARL tasks are often set in static simulated environments. In contrast, LLM-based multi-agent systems are designed for real-life agentic tasks with higher uncertainty. These tasks can be decomposed in various ways, leading to different multi-agent systems with distinct populations and roles. Moreover, agents may exhibit sequential dependencies and contextual relationships when solving sub-tasks. For example, one agent’s action may depend on the outcome of another agent’s action. This dynamic organization can be either human-designed or agent-explored (e.g. , through training), adding another layer of complexity to the design process.
Optimization Space
@@
If we take one LLM generation as an action, the observation space is exponentially vast as $o\in\mathcal{O}\subseteq\mathcal{V}^{L}$, whose complexity is $O(\lvert\mathcal{V}\rvert^{L})$, where $\mathcal{V}$ is the vocabulary of the acting LLM and $N$ is the token length of the generated action, and both of them vary with different acting LLMs. It makes the value function hard to converge with stability. If we take one token as an action, the complexity can be reduced to $O(\lvert\mathcal{V}\rvert\times L)$, but it will become a task with super sparse reward outcome so that the value function should also be meticulously learned, as $L$ can be extremely large. This is also reflected in VAPO and is mitigated by value pretraining (Yue et al., 2025). Fortunately, not all LLM-related tasks face such large optimization spaces. For instance, in embodied AI with LLMs, predefined actions (e.g. , "go to the kitchen," "grab a coffee") are often provided, significantly narrowing the action space (Carta et al., 2023; Tan et al., 2024).
These differences collectively make the optimization of LaMAS more complex than traditional MARL problems. MARFT is designed to tackle the difficulties. To generally optimize the LaMAS regardless of the highly dynamic workflow or organization, MARFT reframes the problem as a sequential decision-making problem despite the dynamic organization. It preserves pretrained utilities by applying clipping, preventing excessive policy drift. Agent profiles are encoded to activate specific capabilities, and no modality assumptions are made, as all agent actions are expressed in tokens from their respective vocabularies.
@@ 4 Multi-Agent Reinforcement Fine-Tuning
Building on the formulations outlined in the previous section, we now turn to the core methodologies that enable MARFT to address the unique challenges of optimizing LaMAS.
@@ 4.1.1 Action-level Fine-Tuning
Action-level MARFT
Inheriting the SM-style re-modeling via the multi-agent advantage decomposition theorem introduced in the last section, a naive variant of MARFT, i.e., Action-level MARFT, is developed and unifies the dynamic LaMAS decision processes perfectly. Its procedure is demonstrated in Figure 3.
@@
Given any organizational or task-solving workflow of a LaMAS, theorem (17) allows MARFT to reframe it as a sequential decision-making process. When collecting interactive trajectories, the LaMAS generates actions with an auto-regressive problem-solving style. The "encoder" constructs local roll-out observations using agent profiles and other relevant information, while the "decoder" combines these observations with dependent predecessors’ actions to generate the current agent’s action $a^{i}\sim p_{\pi^{i}}(a^{i}|\hat{o}^{i},a^{\mathcal{N}(i)})$, where $\mathcal{N}(i)=\left\{j\neq i\mid\mathcal{D}(a^{j},a^{i})=1\right\}$. Then, during training, MARFT adopts an optimization objective in (20). Notably, in naive MARFT, each agent optimizes within a trust region conditioned on predecessor actions, akin to MAT, thus ensuring monotonic improvement as in HAPPO’s sequential update scheme.
$L(\bm{\theta})=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=0}^{T-1}\min\left[\bm{r}_{t}^{i}(\bm{\theta})\hat{A}_{t},\text{clip}\left(\bm{r}_{t}^{i}(\bm{\theta}),1\pm\epsilon\right)\hat{A}_{t}\right],$ (20)
$\text{where}\ \bm{r}_{t}^{i}(\bm{\theta})=\frac{\pi_{\theta^{i}}^{i}\left(a_{t}^{i}|\hat{o}_{t}^{i},\hat{\bm{a}}_{t}^{\mathcal{N}(i)}\right)}{\pi_{\theta^{i}_{\text{old}}}^{i}\left(a_{t}^{i}|\hat{o}_{t}^{i},\hat{\bm{a}}_{t}^{\mathcal{N}(i)}\right)},\;\mathcal{N}(i)=\left\{j\neq i\mid\mathcal{D}(a^{j},a^{i})=1\right\}.$
Action Normalization
@@
As shown in equation (21), the probability of each token $P\left(w|\cdot\right)$ is less than (or equal to) 1. As a result, longer action tends to have a lower joint probability, even though it is more reasonable than other actions in the task. To address the issue, Tan et al. (2024) proposed token normalization and word normalization by dividing the joint action probability in (21) by a constant related to its token or word length. It gets rid of the strong assumption that all the actions have similar lengths, which in LLMs, can be an evil condition, because in agentic tasks, it is all too common that solutions generated can be either super long with a chain-of-thought or super short with a direct answer.
$\text{log}P_{\pi^{i}}\left(a^{i}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)}\right)$ $=\text{log}P_{\pi^{i}}\left(w^{i,1},\ldots,w^{i,L^{i}}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)}\right)$
$=\sum_{j=1}^{L^{i}}\text{log}P_{\pi^{i}}\left(w^{i,j}|\hat{o}^{i},\bm{a}^{\mathcal{N}(i)},w^{i,1},\ldots,w^{i,j-1}\right).$ (21)
Agent-by-agent Update
@@
Updating every policy during each training epoch may introduce implicit off-policy issues, which can potentially compromise the monotonic improvement guarantee of PPO for individual agents’ policies. This problem can be exacerbated in LaMAS, where agents often have dependencies on the results of other agents’ actions. Drawing on the approach of A2PO (Wang et al., 2023), an agent-by-agent update scheme can be employed to restore monotonic improvement for each agent and facilitate stable convergence.
Trajectory-level Optimization
@@
For language models, the definition of action from the RL perspective can be dynamic. Direct trajectory-level optimization, such as GRPO with outcome supervision (Shao et al., 2024) and its variant DAPO (Yu et al., 2025), can be viewed as an action-level optimization with non-zero reward only at the end of the trajectory in the MDP or a normal action-level optimization in a single-step variant of the MDP. Consequently, trajectory-level optimization can be seamlessly integrated into MARFT by following the same idea.
@@ 4.1.2 Token-level Fine-Tuning
Naive Token-level MARFT
On top of the credit assignment of each agent on its action, we can also give a more fine-grained credit assignment by naturally treating every token as an action. In this situation, we need to define the token-level multi-agent observation-action value function and observation value function as
$Q^{\bm{\pi}}\left(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right)}\left[\sum_{k=t}^{\infty}\gamma^{k-t}R_{k}(\bm{o}_{k},\bm{a}_{k})|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1},w_{t}^{i,j}\right],$ (22)
$V^{\bm{\pi}}\left(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)\coloneqq\mathbb{E}_{\tau\sim p_{\bm{\pi}}\left(\tau|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right)}\left[\sum_{k=t}^{\infty}\gamma^{k-t}R_{k}(\bm{o}_{k},\bm{a}_{k})|\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,j-1}\right].$ (23)
And the token-level Bellman backup can be spontaneously derived as
$Q^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j-1},w_{t}^{i,j})\leftarrow\begin{cases}0+\gamma\max\limits_{w_{t}^{i,j+1}}Q^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j},w_{t}^{i,j+1})&\mathrm{if~{}}j<|a_{t}^{i}|\\ 0+\gamma\max\limits_{w_{t}^{i+1,1}}Q^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}{i+1<n}\\ R(\bm{o}_{t},\bm{a}_{t}^{1:i})+\gamma\max\limits_{w_{t+1}^{1,1}}Q^{\bm{\pi}}(\bm{o}_{t+1},\bm{a}_{t}^{1:i},w_{t+1}^{1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}i=n\end{cases},$ (24)
$V^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{i-1},w_{t}^{i,1:j})\leftarrow\begin{cases}0+\gamma V^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{1:i-1},w_{t}^{i,1:j+1})&\mathrm{if~{}}j<|a_{t}^{i}|\\ 0+\gamma V^{\bm{\pi}}(\bm{o}_{t},\bm{a}_{t}^{1:i},w_{t}^{i+1,1})&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}{i<n}\\ R(\bm{o}_{t},\bm{a}_{t}^{1:i})+\gamma V^{\bm{\pi}}(\bm{o}_{t+1},\bm{a}_{t}^{1:i},\emptyset)&\mathrm{if~{}}j=|a_{t}^{i}|~{}\&~{}i=n\end{cases}.$ (25)
This kind of modified token-level Bellman backup guides the computations of TD errors and the advantages of tokens, thus leading to a token-level value training and policy optimization. In the context of the naive token-level MARFT approach, each token is treated as an individual action, thereby introducing a new POMDP formulation, where the reward function is highly sparse and assigns zero rewards to all intermediate tokens and preceding agents, with the reward being allocated solely to the final token of the entire multi-agent system. Although this sparsity in the reward signal poses challenges for learning the value function, it significantly reduces the optimization complexity from $O\left(\lvert\mathcal{V}\rvert^{L}\right)$ to $O\left(\lvert\mathcal{V}\rvert\times L\right)$, as previously discussed in Section 3.3. However, due to the introduction of the different MDP, this naive token-level MARFT optimizes inconsistently with the original optimization problem (Wen et al., 2024b).